randomy A magazine published a study on the ammonia levels near tho exit ramp of a highway tunnel. The data in the table below represent daily ammonia conoentrations (in parts per million) on eight during the aftermoon drive time. Complete parts a through c 1.55 a. Find the range of the ammoria levels. Give the units of measurement for the range The range is 3 ppm 1.52 1.34 1.53 1.64 1.41 140 149 Type an integer or a decimal.) b. Find the variance of the ammonia levels. If possible, give the units of measurement for the variance The varianoe iswith units of ppm squared Round to four decimal p units of ppm squared no units units of ppm
Answer #1

Solution:

Sort the data from smallest to largest

1.34, 1.40,1.41, 1.49, 1.52, 1.53, 1.55, 1.64

So range = (Largest - Smallest) = (1.64-1.34) = 0.3ppm

Solution(b)

Mean = (1.34+1.40+1.41+1.49+1.52+1.53+1.55+1.64)/8 = 11.88/8 = 1.485

Variance = (summation((Xi-mean)^2)/(n-1))

Variance = (1.34-1.485)^2 +(1.40-1.485)^2 +(1.41-1.485)^2+(1.49-1.485)^2+(1.52-1.485)^2+(1.53-1.485)^2+(1.55-1.485)^2 +(1.64-1.485)^2)/(8-1))

Varinace = 0.0093 units of ppm squared.